OpenStudy (anonymous):

Jamie Pruitt is an assistant to the city manager. Her annual salary is $45,000. Next month she will receive a yearly cost-of-living adjustment of 2.5%. In addition, her manager plans to award her a merit increase of 4% for her excellent work during the year. What will Pruitt’s salary be for the coming year?

OpenStudy (princeharryyy):

What do u want answer or solution.

OpenStudy (princeharryyy):

@kayleedilbeck be fast in replying. I don't have much time.

OpenStudy (princeharryyy):

she will receive a yearly cost-of-living adjustment of 2.5% implies rs. 45000*0.025=1125 and from 4% => 45000*0.04=1800 so total 45000+1800+1125=$47925

OpenStudy (anonymous):

Thank you! I was able to figure out another problem with your help too :)

OpenStudy (princeharryyy):

:)